OpenStudy (anonymous):

A long thick cylindrical shell with inner radius a and outer radius b is made of conducting material with permeability (\mu) and carries a slowly varying current I(t) uniformly distributed through the material. Find B , H and E in all regions of space, in the quasi-static approximation.

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

Thanks! Was the quasi-static approximation in saying \[\frac{\partial D}{\partial t} \sim 0\]?

OpenStudy (nikvist):

\[\oint\limits_C\vec{H}\cdot d\vec{l}=\iint\limits_S\left(\vec{J}+\frac{\partial\vec{D}}{\partial t}\right)\cdot d\vec{S}\quad,\quad\vec{J}\gg\frac{\partial\vec{D}}{\partial t}\]

OpenStudy (anonymous):

Thanks!